Please help me solve log2x + logx = 2
Answer by stanbon(75887)

solve log2x + logx = 2
-----
log[2x*x] = 2
---
2x^2 = 10^2
2x^2 = 100
x^2 = 50
--------
x^2 = 5*sqrt(2)
---------------------
Cheers,
Stan H.
====================